Grade 10Physical Chemistry

What volume of 95% sulphuric acid (density=1.85 g/cm^3) and what mass of water must be taken to prepare 100 cm^3 of 15% solution of sulphuric acid(density=1.10 g/cm^3)?

Reba Jesudoss
8 Years agoGrade 10

1 Answer

Rituraj Tiwari
5 Years ago
Total mass of solution
= density × volume
= 1.1 × 100
= 110 g

Mass of 15% H2SO4 in solution
= 110 × 15/100
= 16.5 g

Mass of water in solution
= 110 - 16.5
= 93.5 g

Mass of 95% H2SO4 required
= 100 × 16.5 / 95
= 17.37 g

Volume of 95% H2SO4 required
= 17.37 / 1.85
= 9.4 cm^3
= 9.4 ml

Volume of water required
= 100 - 9.4
= 90.6 ml

Therefore, 9.4 ml of 95% H2SO4 should be added in 90.6 ml water.